# Safeguard Fine-Tuned LLMs Through Pre- and Post-Tuning Model Merging

### v1

@@ Appendix A Domain-Specific Tasks Detail
We select 10,000 samples from the MagiCoder dataset Wei et al. (2024) to improve code generation capabilities. Specifically, we uniformly sampled from each coding languages. We report Pass@10 in our experiment results.
@@
Due the the smaller dataset size of OpenFunctions Patil et al. (2023), we split its full training set into two subsets with ratio $9:1$ for training and validating to enhance API call generation capabilities. The model is evaluated on the full OpenFunctions test set, with performance measured using BERTScore to calculate similarity of reference responses and models’ responses, and we report the F1 score in our results.
@@ B.1 Safety Dataset
Qi et al. (2024) is a dataset that contains 330 harmful instructions for evaluating the harmfulness of LLMs. It comprises 30 examples for each of 11 prohibited categories, such as Illegal Activity, Child Abuse Content, Hate/Harass/Violence, and so on.
@@ B.2 Safety Classifier
$ASR=\frac{R_{yes}}{R_{total}-R_{miss}}$
where $R_{yes}$ is the number of harmful responses classified as "yes", $R_{total}$ is the total number of responses, and $R_{miss}$ is the number of responses that failed to be parsed. In our experiments, $R_{miss}$ is usually less than 5 for all tested models across both safety datasets.
@@ C.1 Prompt Template
For base models, we directly apply their own prompt templates during the training and inference phases. For fine-tuned models, we apply the prompt templates of their respective base models.
@@
The prompt for Gemma-2B-it for the tasks of reasoning, code generation, and tool usage proficiency is shown below:
@@ C.2 Fine-tuning
C.3 Inference
We use greedy decoding to ensure result consistency, except for the HumanEval benchmark. For HumanEval, we apply sampling-based decoding with a temperature of 0.6, top_p of 0.9, top_k of 50, and a repetition penalty of 1.2. To accelerate the inference process, we utilize the VLLM engine for model inference.
@@ D.1 Merging Methods
Linear Merging involves directly combining the weights of the base model and the fine-tuned model by interpolating their parameters. Specifically, the weights of the merged model are calculated as a weighted average of the base and fine-tuned models’ weights, following Equation 3. This method is straightforward and computationally efficient, making it a popular choice for basic model integration.
@@
Spherical Linear Interpolation (SLERP) White (2017) is an advanced merging technique that interpolates between model weights on a hypersphere, ensuring a smoother and more natural transition between the models. Unlike Linear Merging, SLERP accounts for the angular relationship between weight vectors, which aim to better preserve the base model’s features while effectively integrating the fine-tuned model’s task-specific enhancements.
@@
Drop and Rescale (DARE) Yu et al. (2024) is a method used to prepare models for merging techniques such as Linear Merging. It operates by randomly dropping parameters according to a specified drop rate and rescaling the remaining parameters. This process helps reduce the number of redundant and potentially interfering parameters among multiple models. In our experiments, we apply DARE only to the fine-tuned models and then use Equation 3 to merge them with the base model.
Model Stock
Model Stock Jang et al. (2024) utilizes the geometric properties of the weights of fine-tuned models. It determines the optimal merging ratio for each model by minimizing the distance between the merged weights and the center of the weights of the fine-tuned models and the base model. However, since this method requires at least two fine-tuned models and a base model, we choose to merge three models with the base model in our experiments. Therefore, when implementing Model Stock, we first uniformly average the weights of the three fine-tuned models, denoted as
$\theta_{m}=\frac{\theta_{42}+\theta_{1024}+\theta_{48763}}{3},$
where $\theta_{42},\theta_{1024},\theta_{48763}$ represent the model trained with different seed, then we merge this average with the base model using the optimal ratio $\lambda_{optimal}$ determined by Model Stock. The weight of the final merged model is expressed as:
$\theta_{merged}=(1-\lambda_{optimal})\theta_{base}+\lambda_{optimal}\theta_{m}.$
@@ D.2 Model Merging Implementation
For Linear Merging, we determine the interpolation weight $\lambda$ by testing values in the set $\{0.2,0.4,0.6,0.8\}$. For SLERP and DARE, we test the interpolation weight $\lambda$ in the same set, $\{0.2,0.4,0.6,0.8\}$, but do not experiment with the threshold for the dot product in the SLERP algorithm. Regarding Model Stock, since it can automatically approximate the interpolation weight based on the geometric nature of the training trajectory, no hyperparameters need to be specified. We use MergeKit Goddard et al. (2024) as our codebase.
Appendix E Impact of $\lambda$ for Safety
We also illustrate the impact of $\lambda$ on AdvBench in Figure 3. Similar to the results for HEx-PHI, the ASR increases as $\lambda$ increases, but it remains better than the results of the SFT model.
Figure 3: Impact of $\lambda$ on Task Performance and Safety. This figure illustrates how varying $\lambda$ affects the ASR for the AdvBench dataset across different merging methods.

### v2

@@ Appendix A Domain-Specific Tasks Detail
We select 10,000 samples from the MagiCoder dataset Wei et al. (2024) to improve code generation capabilities. Specifically, we uniformly sampled from each coding languages. When evaluating on HumanEval, we set $n=50$, representing the number of responses generated per question, and report Pass@10 as our evaluation metric. During evaluation, we prepend the instruction: "Complete the following code and return only the completed code, without any explanations or additional text." to enforce that the model generates only executable code.
@@
Due to the smaller size of the OpenFunctions dataset Patil et al. (2023), we split its full training set into training and validation subsets using a 9:1 ratio to enhance the model’s API call generation capabilities. The model is evaluated on the full OpenFunctions test set, with performance measured using BERTScore to compute the similarity between the reference responses and the model outputs. We report the F1 score as our evaluation metric. During evaluation, we prepend the instruction: "Complete the following code and return only the completed code, without any explanations or additional text." to ensure that the model generates only executable function calls.
Instruction Following
To assess whether instruction-following ability is preserved after fine-tuning, we evaluate models on IFEval Zhou et al. (2023), a benchmark specifically designed to test instruction adherence. We use the official IFEval evaluation set and report both prompt and instruction accuracy. Since our models are fine-tuned from instruct-tuned variants, this evaluation helps determine whether merging can recover instruction-following capabilities degraded during task-specific fine-tuning. We use lm-evaluation-harness Gao et al. (2024) as the code base for this evaluation.
@@ B.1 Safety Dataset
Qi et al. (2024) is a dataset containing 330 harmful instructions designed to evaluate the harmfulness of LLMs. It includes 30 examples for each of 11 prohibited categories, which are derived from a combination of OpenAI’s usage policies and Meta’s Llama 2 acceptable use policy. These categories are:
  1. 1.
    Illegal Activity
  2. 2.
    Child Abuse Content
  3. 3.
    Hate / Harassment / Violence
  4. 4.
    Malware
  5. 5.
    Physical Harm
  6. 6.
    Economic Harm
  7. 7.
    Fraud / Deception
  8. 8.
    Adult Content
  9. 9.
    Political Campaigning
  10. 10.
    Privacy Violation Activity
  11. 11.
    Tailored Financial Advice
@@ B.2 Safety Classifier
$ASR=\frac{N_{yes}}{N_{total}-N_{miss}}\times 100\%$
where $N_{yes}$ is the number of harmful responses classified as "yes", $N_{total}$ is the total number of responses, and $N_{miss}$ is the number of responses that failed to be parsed. In our experiments, $N_{miss}$ is negligible for all tested models across both safety datasets.
@@ C.1 Prompt Template
For aligned models, we directly apply their own prompt templates during the training and inference phases. For fine-tuned models, we apply the prompt templates of their respective aligned models.
@@
The prompt for Gemma-2 for the tasks of reasoning, code generation, and tool usage proficiency is shown below:
@@
The prompt for Qwen2.5 for the tasks of reasoning, code generation, and tool usage proficiency is shown below:
<|im_start|>system
You are a helpful assistant.
<|im_end|>
<|im_start|>user
{Instruction}
<|im_end|>
<|im_start|>assistant
{Response}
The prompt for the medical assistance task is as follows:
<|im_start|>system
If you are a doctor, please answer the medical
questions based on the patient’s description.
<|im_end|>
<|im_start|>user
{Instruction}
<|im_end|>
<|im_start|>assistant
{Response}
@@ C.2 Fine-tuning
C.3 Baseline Methods
We evaluate dropout rates in the range of $0.1$ to $0.5$, and weight decay coefficients also from $0.1$ to $0.5$. The optimal hyperparameters for each technique are selected based on performance on the downstream tasks validation set.
C.4 Inference
We use greedy decoding to ensure result consistency, except for the HumanEval benchmark. For HumanEval, we apply sampling-based decoding with a temperature of 0.6, top_p of 0.9, top_k of 50, and a repetition penalty of 1.2. To accelerate the inference process, we utilize the vLLM engine Kwon et al. (2023) for model inference.
@@ D.1 Merging Methods
Linear Merging involves directly combining the weights of the aligned model and the fine-tuned model by interpolating their parameters. Specifically, the weights of the merged model are calculated as a weighted average of the base and fine-tuned models’ weights, following Equation 3. This method is straightforward and computationally efficient, making it a popular choice for basic model integration.
@@
Spherical Linear Interpolation (SLERP) White (2017) is an advanced merging technique that interpolates between model weights on a hypersphere, ensuring a smoother and more natural transition between the models. Unlike Linear Merging, SLERP accounts for the angular relationship between weight vectors, which aim to better preserve the aligned model’s features while effectively integrating the fine-tuned model’s task-specific enhancements.
@@
Drop and Rescale (DARE) Yu et al. (2024) is a method used to prepare models for merging techniques such as Linear Merging. It operates by randomly dropping parameters according to a specified drop rate and rescaling the remaining parameters. This process helps reduce the number of redundant and potentially interfering parameters among multiple models.
@@ D.2 Model Merging Implementation
We adopt MergeKit Goddard et al. (2024) as our implementation framework and only vary the interpolation factor $\lambda$. For Linear Merging, we test $\lambda$ values in the range $0.1,0.2,\dots,0.9$} with a step size of 0.1. For SLERP and DARE, we use the same range of $\lambda$ values and follow their respective default configurations in MergeKit—specifically, the default dot product threshold for SLERP and the default drop rate for DARE.
Appendix E More Results
E.1 Comparison of Different Methods
Figure 5: Pareto analysis of downstream task performance and safety across different models and tasks. We present the trade-off between performance and attack success rate (ASR) on HEx-PHI when applying weight decay, dropout, and Linear Merging.
Table 1: Performance and ASR on the downstream task. We compare different merging methods with SFT and baselines. Merging often improves downstream task performance while retaining safety. Bold indicates the best score per metric (excluding Aligned). For the attack success rate on AdvBench and HEx-PHI, we report percentage values. For the downstream tasks of Reasoning and Code Generation, we report accuracy, while for the remaining two tasks, we report F1 scores.
Task Method LLaMA-3-8B-Instruct Gemma-2-2B-It Qwen2.5-7B-Instruct Perf. $\uparrow$ AdvBench $\downarrow$ HEx-PHI $\downarrow$ Perf. $\uparrow$ AdvBench $\downarrow$ HEx-PHI $\downarrow$ Perf. $\uparrow$ AdvBench $\downarrow$ HEx-PHI $\downarrow$ Reasoning Aligned 61.30% 0.00% 1.22% 28.98% 0.58% 0.00% 24.16% 0.38% 9.09% SFT 67.84% 4.25% 12.41% 39.16% 0.38% 0.51% 65.94% 2.44% 17.88% Weight Decay 67.85% 15.38% 30.71% 39.41% 0.19% 0.71% 65.92% 3.78% 20.40% Dropout 67.83% 16.79% 35.96% 39.89% 0.96% 0.71% 66.45% 4.49% 24.55% Linear 69.23% 0.64% 6.38% 40.07% 0.06% 0.00% 66.96% 1.03% 12.32% DARE 68.64% 1.28% 5.66% 40.01% 0.10% 0.00% 66.89% 1.09% 12.22% SLERP 68.68% 1.22% 5.86% 40.05% 0.26% 0.00% 66.73% 0.96% 13.03% Medical Assistance Aligned 0.5242 0.00% 1.22% 0.5151 0.58% 0.00% 0.5271 0.38% 9.09% SFT 0.5711 30.06% 38.85% 0.5254 1.41% 1.92% 0.5751 0.77% 12.42% Weight Decay 0.5740 23.33% 32.22% 0.5594 2.37% 7.47% 0.5631 0.58% 8.28% Dropout 0.5744 22.31% 31.41% 0.5632 3.59% 7.07% 0.5226 0.71% 7.68% Linear 0.5738 0.32% 4.06% 0.5243 1.15% 1.21% 0.5721 0.45% 11.11% DARE 0.5758 5.61% 23.41% 0.5248 1.15% 1.21% 0.5724 0.26% 11.52% SLERP 0.5789 5.76% 24.26% 0.5243 1.15% 1.52% 0.5729 0.32% 11.72% Code Generation Aligned 71.63% 0.00% 1.22% 51.96% 0.58% 0.00% 85.89% 0.38% 9.09% SFT 74.19% 2.25% 11.67% 52.63% 2.76% 5.76% 88.06% 0.64% 7.98% Weight Decay 73.47% 1.67% 8.08% 53.20% 2.44% 6.97% 88.08% 0.71% 13.74% Dropout 73.64% 1.73% 8.23% 53.17% 2.95% 5.96% 87.70% 0.83% 11.52% Linear 75.32% 0.71% 4.27% 53.04% 1.73% 3.03% 89.37% 0.32% 7.88% DARE 74.46% 0.64% 4.65% 53.09% 1.86% 3.74% 89.64% 0.51% 7.07% SLERP 75.01% 0.71% 4.34% 53.07% 1.67% 3.23% 89.39% 0.32% 8.18% Tool Using Proficiency Aligned 0.8979% 0.00% 1.22% 0.7280 0.58% 0.00% 0.9357 0.38% 9.09% SFT 0.8989 0.83% 3.45% 0.8802 0.64% 0.10% 0.9369 0.58% 8.08% Weight Decay 0.9282 1.41% 3.22% 0.8838 0.77% 0.30% 0.9177 0.58% 8.48% Dropout 0.9269 0.83% 1.92% 0.8865 0.83% 0.40% 0.9514 0.77% 10.90% Linear 0.9266 0.77% 2.44% 0.8793 0.64% 0.20% 0.9489 0.13% 9.39% DARE 0.9251 0.45% 1.21% 0.8793 0.64% 0.20% 0.9149 0.06% 9.39% SLERP 0.9266 0.44% 1.72% 0.8802 0.64% 0.10% 0.9152 0.13% 9.19%
In Section 5.1, we demonstrate that Linear Merging consistently achieves a better trade-off between performance and safety when evaluated on various downstream tasks and AdvBench. Figure 5 further confirms this trend on the HEx-PHI benchmark, where Linear Merging yields favorable Pareto fronts across different models and tasks.
To better reflect practical usage scenarios, we additionally report results based on the best-performing model (on the validation set of each task) within each method category—including Weight Decay, Dropout, Linear Merging, DARE, and SLERP. These results are summarized in Table 1, providing a fair comparison of each method’s effectiveness under optimal conditions. We use validation set performance for model selection, as it is commonly available during deployment and serves as a realistic basis for method comparison.
In Table 1, even when each method is allowed to select its best-performing checkpoint, merging-based approaches still exhibit strong capability in recovering the safety of the fine-tuned model, often outperforming regularization-based methods such as Dropout and Weight Decay. This suggests that model merging is not only effective but also practical for mitigating safety degradation in real-world settings, even without access to additional safety data.
E.2 Which safety category suffers the most from safety degradation?
In this section, we investigate which categories in HEx-PHI are most affected by safety degradation. All categories are listed in Appendix B.1.
As observed in Section 5.1, LLaMA-3-8B-Instruct and Qwen2.5-7B-Instruct exhibit the most severe degradation on the Reasoning and Medical Assistance tasks. Therefore, we analyze their responses on the HEx-PHI benchmark to further understand which safety categories are most impacted.
The category distributions are shown in Figure 6. For LLaMA-3-8B-Instruct, the aligned model only generates harmful responses in categories 4 (Malware), 9 (Political Campaigning), and 10 (Privacy Violation Activity). After fine-tuning, however, harmful responses increase across all categories, with categories 4, 7 (Fraud/Deception), and 9 exhibiting the most significant growth in both tasks. This demonstrates that safety degradation extends to fine-grained category levels, making it difficult to address safety concerns solely by modifying the model prior to fine-tuning.
Qwen2.5-7B-Instruct shows a slightly different trend. Its aligned model generates harmful responses across more categories compared to LLaMA-3-8B-Instruct, and fine-tuning further aggravates these issues. However, both models share a common pattern: a large number of harmful responses appear in categories 7 and 9. This suggests that certain categories may be particularly vulnerable to safety degradation during fine-tuning, regardless of model architecture or downstream task.
After applying different merging methods, most harmful categories show a reduction in the number of harmful responses. However, the degree of improvement varies across merging strategies and tasks. For instance, Linear Merging performs best on LLaMA-3-8B-Instruct but not on Qwen2.5-7B-Instruct, and some categories do not benefit from merging at all. This indicates that no single method universally outperforms others in preserving safety across all harmful categories.
LLaMA-3-8B-Instruct (Reasoning)
LLaMA-3-8B-Instruct (Medical Assistance)
Qwen2.5-7B-Instruct (Reasoning)
Qwen2.5-7B-Instruct (Medical Assistance)
Figure 6: Safety degradation across categories in the HEx-PHI benchmark. ASR distributions over 11 harmful categories for LLaMA-3-8B-Instruct and Qwen2.5-7B-Instruct on the Reasoning and Medical Assistance tasks.